\begin{thebibliography}{99}


\bibitem{At-Bo} M.F. Atiyah, R. Bott: A Lefschetz fixed point
formula for elliptic complexes, I. Ann. Math, 86, 374-407, (1967).

\bibitem{At-Si} M.F. Atiyah, I.M. Singer: The index of elliptic
operators, I, Ann. Math., 87, 484-530, (1968).

\bibitem{Az1} R. Azencott: Formule de Taylor stochastique et
d\'eveloppements asymptotiques d'int\'egrales de Feynman. In
Azema, Yor (Eds.), S\'eminaire de probabilit\'es \textbf{XVI}, LNM
\textbf{921}, 237-284, Springer, (1982).


\bibitem{Bau} F. Baudoin: An Introduction to the Geometry of
Stochastic Flows, Imperial College Press, (2004).

%\bibitem{Bau2} F. Baudoin: The tangent space to a hypoelliptic diffusion
%and applications, S\'eminaire de Probabilit\'es \textbf{XXXVIII},
%(2004) .


%\bibitem{Bel} Bella\"iche A.: The tangent space in sub-Riemannian
%geometry, In Sub-Riemannian Geometry, edited by A. Bella\"iche and
%J.J. Risler, Birkh\"auser, (1996).

\bibitem{Ben} G. Ben Arous: D\'eveloppement asymptotique du noyau de la chaleur hypoelliptique sur la diagonale,
Annales de l'institut Fourier, tome \textbf{39}, p. 73-99, (1989).

\bibitem{Ben2} G. Ben Arous: Flots et s\'eries de Taylor
stochastiques, Journal of Probability Theory and Related Fields,
\textbf{81}, 29-77, (1989).

\bibitem{Ber-Ge-Ve} N. Berline, E. Getzler, M. Vergne: Heat
kernels and Dirac operators, Grundlehren Text Editions, Second
edition, Springer, (2003).

\bibitem{Bi2}  J.M. Bismut: The Atiyah-Singer Theorems: A Probabilistic
Approach, J. of Func. Anal., Part I: \textbf{57} (1984), Part II:
\textbf{57}, 329-348, (1984).

%\bibitem{Bour} N. Bourbaki: Groupes et Alg\`ebres de Lie, Chap
%1-3, Hermann, (1972).

\bibitem{Cast} F. Castell: Asymptotic expansion of stochastic
flows, Prob. Rel. Fields, \textbf{96}, 225-239, (1993).

\bibitem{Che} K.T. Chen: Integration of paths, Geometric invariants and a Generalized Baker-Hausdorff formula,
Annals of Mathematics, \textbf{65}, n°1, (1957).

%\bibitem{Chern} S.S. Chern: A simple intrinsic proof of the Gauss-Bonnet theorem
%for closed Riemannian manifolds, Ann. Math. 45, 747-752, (1944).


\bibitem{Fli} M. Fliess, D. Normand-Cyrot: Alg\`ebres de Lie
nilpotentes, formule de Baker-Campbell-Hausdorff et int\'egrales
it\'er\'ees de K.T. Chen, in S\'eminaire de Probabilit\'es, LNM
\textbf{920}, Springer-Verlag, (1982).



%\bibitem{Ga} B. Gaveau: Principe de moindre action, propagation de la chaleur et estim\'ees sous-elliptiques sur
%certains groupes nilpotents. Acta Math. 139 (1-2), 95-153, (1977).

\bibitem{Get} E. Getzler: A short proof of the Atiyah-Singer index
theorem. Topology, 25, 111-117, (1986).

\bibitem{Gi} P.B. Gilkey: Curvature and the eigenvalues of the
Laplacian for elliptic complexes, Advances in Math., 10, 344-382,
(1973).

\bibitem{Hsu} E.P. Hsu: Stochastic Analysis on manifolds, AMS, Graduate Texts in Mathematics, Volume \textbf{38},
(2002).



\bibitem{Leandre} R. L\'eandre: Sur le th\'eor\`eme d'Atiyah-Singer, Probab. Theory Related Fields  80,  no. 1,
119-137, (1988).

%\bibitem{Leandre2} R. L\'eandre: D\'eveloppement asymptotique de la
%densit\'e d'une diffusion d\'eg\'en\'er\'ee, Forum Math.
%\textbf{4}, \textbf{1}, 45-75, (1992).


\bibitem{Ly} T. Lyons: Differential Equations Driven by Rough Signals.Revista Mathem\`atica
Iberio Americana, Vol \textbf{14}, No \textbf{2}, 215 - 310,
(1998).

\bibitem{Ly-Vi} T. Lyons, N. Victoir: Cubature on Wiener space, Proc. R. Soc. Lond. A
(2004) 460, 169-198.

\bibitem{McK-Sin} H. McKean, I.M. Singer: Curvature and the
eigenvalues of the Laplacian. J. Diff. Geometry, 1, 43-69, (1967).

\bibitem{Nu} Nualart D.: The Malliavin calculus and related topics,
Springer, Berlin Heidelberg New-York (1995).

%\bibitem{Mal1} P. Malliavin: Stochastic calculus of variations and hypoelliptic operators. In: \textit{Proc.
%Inter, Symp. on Stoch. Diff. Equations, Kyoto 1976},  Wiley,
%195-263, (1978).

\bibitem{Mal2} P. Malliavin: Stochastic Analysis, Grundlehren der
mathematischen Wissenschaften, Vol. \textbf{313}, Springer,
(1997).



\bibitem{Pat} V.K. Patodi: An analytic proof of the Riemann-Roch
-Hirzebruch theorem, J. Diff. Geom. 5, 251-283, (1971).

%\bibitem{Reu} C. Reutenauer: Free Lie algebras, London
%Mathematical Society Monographs, New series \textbf{7}, (1993).

%\bibitem{Ros} S. Rosenberg: The Laplacian on a Riemannian manifold, London Mathematical Society Student Texts, 31, (1997).


%\bibitem{Ro-St} L.P. Rotschild, E.M. Stein: Hypoelliptic
%differential operators and Nilpotent Groups, Acta Mathematica,
%\textbf{137}, 247-320, (1976).

\bibitem{Stri} R.S. Strichartz: The Campbell-Baker-Hausdorff-Dynkin formula and solutions of
differential equations, Jour. Func. Anal., \textbf{72}, 320-345,
(1987).

%\bibitem{Tak} S. Takanobu (1988): Diagonal short time asymptotics of heat
%kernels for certain degenerate second order differential operators
%of H\"ormander type. Publ. Res. Inst. Math. Sci \textbf{24},
%169-203, (1988).

%\bibitem{Tay2} M.E. Taylor: Partial Differential Equations,
%Qualitative Studies of Linear Equations, Springer, Applied
%Mathematical Sciences \textbf{116}, (1996).

%\bibitem{Yor} M. Yor: Exponential functionals of Brownian motion and Related processes, Springer Finance, (2001).



\end{thebibliography}

\documentclass[11pt]{article}

\usepackage{amsfonts}

\usepackage{amsmath}

\def\theequation{\thesection.\arabic{equation}}

\newcommand{\beq}{\begin{equation}}

\newcommand{\eeq}{\end{equation}}

\newcommand{\bea}{\begin{eqnarray}}

\newcommand{\eea}{\end{eqnarray}}

\newcommand{\beas}{\begin{eqnarray*}}

\newcommand{\eeas}{\end{eqnarray*}}

\newcommand{\ds}{\displaystyle}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{remark}[theorem]{Remark}

\newtheorem{example}[theorem]{Example}

\newtheorem{examples}[theorem]{Examples}

\newtheorem{foo}[theorem]{Remarks}

\newenvironment{proof}{\addvspace{\medskipamount}\par\noindent{\it

Proof}.}{\unskip\nobreak\hfill$\Box$\par\addvspace{\medskipamount}}

\newcommand{\ang}[1]{\left<#1\right>}  % angular brackets for projection

\newcommand{\brak}[1]{\left(#1\right)}    % round brackets

\newcommand{\crl}[1]{\left\{#1\right\}}   % curly brackets

\newcommand{\edg}[1]{\left[#1\right]}     % edgy brackets

\newcommand{\E}[1]{{\rm E}\left[#1\right]}

\newcommand{\var}[1]{{\rm Var}\left(#1\right)}

\newcommand{\cov}[2]{{\rm Cov}\left(#1,#2\right)}

\newcommand{\N}[1]{||#1||}     % Norm

\newcommand{\abs}[1]{\left|#1\right|}     % absolute value

\title{Chen series and Atiyah-Singer theorem}

\author{Fabrice Baudoin
\\
\\\small Laboratoire de Statistiques et Probabilit\'es
\\\small Universit\'e Paul Sabatier
\\\small 118 Route de Narbonne, Toulouse, France
\\\small fbaudoin@cict.fr}

\date{}

\begin{document}

\maketitle

\begin{abstract}
The purpose of this work is to give a new and short proof of the
Atiyah-Singer local index theorem for the Dirac operator on the
spin bundle. This proof is obtained by using heat semigroups
approximations
 based on the truncation of Brownian Chen series.

\end{abstract}

\tableofcontents

\section{Introduction}

\

The goal of this paper is to give a new and short proof of the
local Atiyah-Singer index theorem by using approximations of heat
semigroups. The heat equation approach to index theorems is not
new: It was suggested by Atiyah-Bott \cite{At-Bo} and
McKean-Singer \cite{McK-Sin}, and first carried out by Patodi
\cite{Pat} and Gilkey \cite{Gi}. Bismut in \cite{Bi2} introduces
stochastic methods based on Feynman-Kac formula. For probabilistic
approaches that are mainly based on Bismut's ideas, we also refer
to \cite{Leandre} and Chapter 7 of \cite{Hsu}. For a complete
survey on (non probabilistic) heat equation methods for index
theorems, we refer to the book \cite{Ber-Ge-Ve}.

\

However, in our approach we will see that the $A$-genus appears in
a natural way,  from purely local computations on approximations
of  heat semigroups. Our method relies on explicit approximations
of the holonomy over the heat equation on vector bundles and
unlike the other probabilistic approaches does not involve the
Feynman-Kac formula nor the technics of stochastic differential
geometry.

\

The idea is the following. Let $\mathbf{P}_t$ denote a heat
semigroup. In recent works, see Baudoin \cite{Bau} and
Lyons-Victoir \cite{Ly-Vi}, by using Brownian Chen series, it has
been pointed out that $\mathbf{P}_t$ admits a formal
representation as the expectation of the exponential of a random
Lie series. The truncature of this Lie series leads to explicit
approximations of $\mathbf{P}_t$. More precisely, one gets a
family of operators $\mathbf{P}^N_t$, $N \ge 1$, such that (in the
$\sup$ norm)
\begin{align}\label{pr}
\mathbf{P}_t=\mathbf{P}_t^N + O(t^{\frac{N+1}{2}}), \quad t
\rightarrow 0.
\end{align}
This point of view has  been used in Lyons-Victoir \cite{Ly-Vi} to
provide cubature formulae on Wiener space that give efficient
numerical approximations of solutions of heat equations.

\

Assume now that $\mathbf{P}_t$ is the heat semigroup associated
with the Dirac operator on the Clifford module over a compact
$d$-dimensional spin manifold, $d$ even. From (\ref{pr}), we will
classically deduce
\[
\mathbf{Str} \text{ } \mathbf{P}_t=\mathbf{Str} \text{ }
\mathbf{P}_t^d + O(t^{\frac{1}{2}}), \quad t \rightarrow 0,
\]
where $\mathbf{Str}$ denotes the supertrace. The Lie structure
that explicitly appears in $\mathbf{P}_t^d$ now leads to algebraic
cancellations that imply
\[
\mathbf{Str} \text{ } \mathbf{P}^d_t=\mathbf{Str} \text{ }
\mathbf{P}_t^2 + O(t^{\frac{1}{2}}), \quad t \rightarrow 0.
\]
Since, from McKean-Singer theorem, the supertrace of
$\mathbf{P}_t$ has to be constant and equal to the index of the
Dirac operator, the local index theorem follows from the easy
computation of $\mathbf{Str} \text{ } \mathbf{P}_t^2$.

\

The paper is organized as follows. In the first part,  we survey
results on random Chen series that are needed for the construction
of approximations of heat semigroups. In the second part, we use
these series to construct explicit approximations of the holonomy
above general heat equations on vector bundles.  Finally, in the
third part, we develop the idea hinted above to provide a new
short proof of the Atiyah-Singer local index theorem for the Dirac
operator on the spin bundle.

\

\textbf{Acknowledgment:} \textit{The author would like to thank M.
Hairer for stimulating discussions}.

\section{Chen series}

We introduce here notations that will be used throughout the paper
and survey some results on Chen series that will be later needed.
Basic background on Chen series with respect to regular paths can
be found in \cite{Che} and background on Chen series with respect
to Brownian paths can be found in Chapter 1 of \cite{Bau} (see
also \cite{Fli} and \cite{Ly-Vi}). Let us note that the Chen
series with respect to Brownian paths is also called, in the rough
paths theory of Lyons, the signature of the Brownian motion.

Let $\mathbb{R} [[X_0,...,X_d]]$ be the non commutative algebra
over $\mathbb{R}$ of the formal series with $d+1$ indeterminates,
that is the set of series
\[
Y=\sum_{k \geq 0} \sum_{i_1,...,i_k} a_{i_1,...,i_k}
X_{i_1}...X_{i_k}.
\]

The exponential of $Y \in \mathbb{R} [[ X_0 ,..., X_d ]]$ is
defined by
\[
\exp (Y)=\sum_{k=0}^{+\infty} \frac{Y^k}{k!}.
\]
We define the bracket between two elements $U$ and $V$ of
$\mathbb{R} [[ X_0 ,..., X_d ]]$ by
\[
[U,V]=UV-VU.
\]
If $I=(i_1,...,i_k) \in \{ 0,..., d \}^k$ is a word, we denote by
$X_I$ the commutator defined by
\[
X_I = [X_{i_1},[X_{i_2},...,[X_{i_{k-1}}, X_{i_{k}}]...].
\]
We will denote by $\mathcal{S}_k$ the set of the permutations of
$\{0,...,k\}$.  If $\sigma \in \mathcal{S}_k$, we denote
$e(\sigma)$ the cardinality of the set
\[
\{ j \in \{0,...,k-1 \} , \sigma (j) > \sigma(j+1) \},
\]
and $\sigma (I)$ the word $(i_{\sigma(1)},...,i_{\sigma(k)})$.

Let us now consider a $d$-dimensional standard Brownian motion
$(B_t)_{t \geq 0}=(B^1_t,...,B^d)_{t \geq 0}$. We use the
convention that $B^0_t=t$. If $I=(i_1,...i_k) \in \{0,...,d\}^k$
is a word with length $k$, the iterated Stratonovich integral
\begin{equation*}
\int_{\Delta^k [0,t]} \circ dB^I= \int_{0 \leq t_1 \leq ... \leq
t_k \leq t} \circ dB^{i_1}_{t_1} \circ ... \circ dB^{i_k}_{t_k},
\end{equation*}
can be defined as the limit in $p$-variation, $p>2$,
\[
\lim_{ n \rightarrow +\infty} \int_{\Delta^k [0,t]}  (dB^n)^I,
\]
where $B^n$ denotes the piecewise linear interpolation of the
paths of  $(B_u)_{0 \le u \le t}$ along the dyadic subdivision of
$[0,t]$.

With the notation,
\[
\Lambda_I (B)_t= \sum_{\sigma \in \mathcal{S}_k} \frac{\left(
-1\right) ^{e(\sigma )}}{k^{2}\left(
\begin{array}{l}
k-1 \\
e(\sigma )
\end{array}
\right) } \int_{\Delta^k [0,t]} \circ dB^{\sigma^{-1}(I)},
\]
we have the following theorem:

\begin{theorem}\label{Chen-Strichartz brownien}
\[
1 + \sum_{k=1}^{+\infty} \sum_{I \in \{0,1,...,d\}^k} \left(
\int_{\Delta^k [0,t]} \circ  dB^I \right) X_{i_1} ... X_{i_k}
=\exp \left( \sum_{k \geq 1} \sum_{I \in \{0,1,...,d\}^k}\Lambda_I
(B)_t X_I \right), \text{ }t \geq 0.
\]
\end{theorem}
This theorem is due to Chen \cite{Che} and Strichartz \cite{Stri}
that prove that the above result holds for absolutely continuous
paths. The result for Brownian paths is pointed out in Fliess
\cite{Fli}. And finally, Lyons \cite{Ly}, with rough paths theory,
shows that it actually can be extended to very general paths.

If
\[
Y=\sum_{k \geq 0} \sum_{I=(i_1,...,i_k)} a_{i_1,...,i_k}
X_{i_1}...X_{i_k}
\]
is a  random series, that is if the coefficients are real random
variables defined on a probability space, we will denote
\[
\mathbb{E}(Y)=\sum_{k \geq 0} \sum_{I=(i_1,...,i_k)}
\mathbb{E}(a_{i_1,...,i_k}) X_{i_1}...X_{i_k}
\]
as soon as this expression makes sense, that is as soon as for
every $I=(i_1,...,i_k)$,
\[
\mathbb{E}(\mid a_{i_1,...,i_k} \mid)<+\infty,
\]
where $\mathbb{E}$ stands for the expectation.

\begin{theorem} \label{main theorem} (See \cite{Bau},
\cite{Ly-Vi}). We have
\[
\exp \left( t \left(X_0+\frac{1}{2}\sum_{i=1}^d X_i^2 \right)
\right)=\mathbb{E} \left( \exp \left( \sum_{k \geq 1} \sum_{I \in
\{0,1,...,d\}^k} \Lambda_I (B)_t X_I \right) \right), \quad t \ge
0.
\]
\end{theorem}

\section{Approximation of elliptic  heat kernels on vector bundles}

In the spirit of Azencott \cite{Az1}, Ben Arous \cite{Ben2},
Castell \cite{Cast} and Lyons-Victoir \cite{Ly-Vi}, we use in this
section Brownian Chen series in order to provide efficient
approximations of heat semigroups and corresponding heat kernels.
The idea is to truncate the Lie series that appear in the  formal
representation of the heat semigroup given by Theorem \ref{main
theorem}. As we shall see, this truncation that has already found
applications for cubature formulae \cite{Ly-Vi}, is also
particularly efficient to approximate the holonomy over the heat
equation in a vector bundle.

\

Let $\mathbb{M}$ be a $d$-dimensional compact smooth Riemannian
manifold and let $\mathcal{E}$ be a finite-dimensional vector
bundle over $\mathbb{M}$. We denote by $\Gamma ( \mathbb{M},
\mathcal{E} )$ the space of smooth sections. Let now $\nabla$
denote a connection on $\mathcal{E}$.

We consider the following linear partial differential equation
\begin{equation}\label{Hormander PDE}
\frac{\partial \Phi}{\partial t}=\mathcal{L} \Phi,\quad
\Phi(0,x)=f(x),
\end{equation}
where $\mathcal{L}$ is an operator on $\mathcal{E}$  that can be
written
\[
\mathcal{L}=\nabla_{0}+\frac{1}{2} \sum_{i=1}^d \nabla_{i}^2,
\]
with
\[
\nabla_i=\mathcal{F}_i +\nabla_{V_i}, \quad 0 \le i \le d,
\]
the $V_i$'s being smooth vector fields on $\mathbb{M}$ and the
$\mathcal{F}_i$'s being smooth potentials (that is sections of the
bundle $\mathbf{End}(\mathcal{E})$). It is known that the solution
of (\ref{Hormander PDE}) can be written
\[
\Phi (t,x)=(e^{t \mathcal{L}}f)(x)=\mathbf{P}_t f (x).
\]
If $I\in \{0,1,...,d\}^k$ is a word, we denote
\[
\nabla_I= [\nabla_{i_1},[\nabla_{i_2},...,[\nabla_{i_{k-1}},
\nabla_{i_{k}}]...].
\]
and
\[
d(I)=k+n(I),
\]
where $n(I)$ is the number of 0 in the word $I$.

For $N \ge 1$, let us consider
\[
\mathbf{P}^N_t=\mathbb{E} \left( \exp \left( \sum_{I, d(I) \le N }
\Lambda_I (B)_t \nabla_{I} \right) \right).
\]

For instance
\[
\mathbf{P}^1_t=\mathbb{E} \left( \exp \left(\sum_{i=1}^d B^i_t
\nabla_{i} \right) \right),
\]
and
\[
\mathbf{P}^2_t=\mathbb{E} \left( \exp \left( \sum_{i=0}^d B^i_t
\nabla_{i} +\frac{1}{2}\sum_{1\leq i < j \leq d} \int_0^t B^i_s
dB^j_s-B^j_s dB^i_s [\nabla_{i},\nabla_{j}] \right) \right).
\]

The meaning of this last notation is the following. If $f \in
\Gamma ( \mathbb{M}, \mathcal{E} )$, then $ (\mathbf{P}^N_t
f)(x)=\mathbb{E} (\Psi (1 ,x))$, where $\Psi (\tau , x)$ is the
solution of the first order partial differential equation with
random coefficients:
\[
\frac{\partial \Psi}{\partial \tau}(\tau ,x)=\sum_{I, d(I) \le N }
\Lambda_I (B)_t (\nabla_{I} \Psi) (\tau ,x), \quad \Psi
(0,x)=f(x).
\]

Let us consider the following family of norms: If $f \in \Gamma (
\mathbb{M}, \mathcal{E} ) $, for $k \ge 0$,
\[
\parallel f \parallel_k =\sup_{0 \le l \le k} \sup_{0\le i_1, \cdots, i_l \le d}
\sup_{x \in \mathbb{M}} \parallel \nabla_{i_1} \cdots \nabla_{i_l}
f (x) \parallel.
\]
%We will use the same notation for sections on $\mathbb{M} \times
%\mathbb{M}$.

\begin{theorem}\label{Parametrix} Let $N \ge 1$ and $k \ge 0$.
For $f \in \Gamma ( \mathbb{M}, \mathcal{E} ) $,
\[
\parallel \mathbf{P}_t f -\mathbf{P}_t^N f \parallel_k=O\left( t^{\frac{N+1}{2}} \right),
\quad t \rightarrow 0.
\]
\end{theorem}

\begin{proof}

First, by using the scaling property of Brownian motion and
expanding out the exponential with Taylor formula we obtain
\[
\exp \left( \sum_{I, d(I) \le N } \Lambda_I (B)_t \nabla_{I}
\right)f=\left( \sum_{k=0}^N \frac{1}{k!} \left(\sum_{I, d(I) \le
N } \Lambda_I (B)_t \nabla_{I}\right)^k\right)f +t^{\frac{N+1}{2}}
\mathbf{R}^1_N (t),
\]
where the remainder term $\mathbf{R}^1_N (t)$ is such that
$\mathbb{E} \left( \parallel \mathbf{R}^1_N (t) \parallel_k
\right)$ is bounded when $t \rightarrow 0$. We now observe that,
due to Theorem \ref{Chen-Strichartz brownien}, the rearrangement
of terms in the previous formula gives

\[
\left( \sum_{k=0}^N \frac{1}{k!} \left(\sum_{I, d(I) \le N }
\Lambda_I (B)_t \nabla_{I}\right)^k\right)f=f + \sum_{I, d(I) \le
N}\int_{\Delta^{\mid I \mid} [0,t]}  \circ dB^I \nabla_{i_1}
...\nabla_{i_{\mid I \mid}}f +t^{\frac{N+1}{2}} \mathbf{R}^2_N
(t),
\]
where $\mathbb{E} \left( \parallel \mathbf{R}^2_N (t) \parallel_k
\right)$ is bounded when $t \rightarrow 0$. Therefore
\[
\exp \left( \sum_{I, d(I) \le N } \Lambda_I (B)_t \nabla_{I}
\right)f=f + \sum_{I, d(I) \le N}\int_{\Delta^{\mid I \mid} [0,t]}
\circ dB^I \nabla_{i_1} ...\nabla_{i_{\mid I \mid}}f
+t^{\frac{N+1}{2}} \mathbf{R}^3_N (t),
\]
and
\[
\mathbf{P}_t^N f=f + \sum_{I, d(I) \le N}\mathbb{E}\left(
\int_{\Delta^{\mid I \mid} [0,t]} \circ dB^I \right) \nabla_{i_1}
...\nabla_{i_{\mid I \mid}}f +t^{\frac{N+1}{2}}\mathbb{E} \left(
\mathbf{R}^3_N (t) \right),
\]
where $\mathbb{E} \left( \parallel \mathbf{R}^3_N (t)
\parallel_k \right)$ is bounded when $t \rightarrow 0$. We now
have to compute the expectation of iterated Stratonovitch
integrals. An easy computation shows that if $\mathcal{I}_n$ is
the set of words  with length $n$ obtained by all the possible
concatenations of the words
\[
\{ 0 \}, \{ (i,i) \}, \quad i \in \{1,...,d\},
\]
\begin{enumerate}
\item If $I \notin \mathcal{I}_n$ then
\[
\mathbb{E} \left( \int_{\Delta^n [0,t]}  \circ dB^I \right) =0 ;
\]
\item If $I \in \mathcal{I}_n$ then
\[
\mathbb{E} \left( \int_{\Delta^n [0,t]}  \circ dB^I \right)
=\frac{t^{\frac{n+n(I)}{2}}}{2^{\frac{n-n(I)}{2}}\left(\frac{n+n(I)}{2}
\right) ! },
\]
where $n(I)$ is the number of 0 in $I$ (observe that since $I \in
\mathcal{I}_n$, $n$ and $n(I)$ necessarily have the same parity).
\end{enumerate}
We conclude therefore
\[
\parallel \mathbf{P}^N_t f -\sum_{k  \le \frac{N+1}{2}} \frac{t^k}{k!} \mathcal{L}^k f\parallel_k =
O\left( t^{\frac{N+1}{2}} \right).
\]
Since it is known that
\[
\parallel \mathbf{P}_t f -\sum_{k  \le \frac{N+1}{2}} \frac{t^k}{k!} \mathcal{L}^k f \parallel_k =
O\left( t^{\frac{N+1}{2}} \right),
\]
the theorem is proved.

\end{proof}

Let us, here, assume that the operator $\mathcal{L}$ is elliptic
at $x_0 \in \mathbb{M}$ in the sense that $(V_1 (x_0),...,V_d
(x_0))$ is an orthonormal basis of the tangent space at $x_0$. In that case,
$\mathbf{P}_t$ is known to admit a smooth Schwartz kernel at
$x_0$. That is, there exists a smooth map
\begin{align*}
p(x_0,\cdot): \mathbb{R}_{>0}\rightarrow \Gamma( \mathbb{M},
\mathbf{Hom} (\mathcal{E}))
\end{align*}
\begin{theorem}
Let $N \ge 1$. There exists a map 
\begin{align*}
p^N(x_0,\cdot): \mathbb{R}_{>0}\rightarrow \Gamma( \mathbb{M},
\mathbf{Hom} (\mathcal{E}))
\end{align*}
such that for $f \in \Gamma ( \mathbb{M}, \mathcal{E} ) $,
\[
(\mathbf{P}^N_t f)(x_0)=\int_\mathbb{M} p^N_t (x_0,y)f(y)dy.
\]
Moreover,
\[
p_t (x_0,x_0)=p_t^N (x_0,x_0)+0\left( t^{\frac{N+1-d}{2}} \right).
\]
\end{theorem}

\begin{proof}

The proof is not simple. We shall proceed in several steps. In a
first step, we shall show the existence of a  kernel at $x_0$ for
$\mathbf{P}^N_t$ acting on functions. In a second step we shall
deduce by parallel transport, the existence of $p^N (x_0,\cdot)$.
And finally, we shall prove the required estimate.

\

\textbf{First step:}

Let us define,
\[
\mathbf{Q}_t^N=\mathbb{E} \left( \exp \left( \sum_{I, d(I) \le N }
\Lambda_I (B)_t V_{I} \right) \right).
\]
In order to show  that $\mathbf{Q}_t^N$ admits a kernel at $x_0$, we show that
the stochastic process
\[
Z_t^N=\exp \left( \sum_{I, d(I) \le N } \Lambda_I (B)_t V_{I} \right) (x_0)
\]
has a density with respect to the Riemannian measure of $\mathbb{M}$. To this end, from

Though the family $(\mathbf{Q}_t^N)_{t \ge 0}$ is not a semigroup
in general, the idea is to factorize $(\mathbf{Q}_t^N)_{t \ge 0}$
via a semigroup acting on functions  and defined on a $N$-step
nilpotent group.

\

Up to isomorphism, there exists a unique simply connected and
nilpotent Lie group $\mathbb{G}_N$ with Lie algebra
$\mathfrak{g}_N $ such that  $\mathfrak{g}_N $ is isomorphic by a
Lie algebra morphism $\Phi$ to the Lie algebra $\mathbb{R}
[[X_0,...,X_d]]$ quotiented by the relations
\[
\{ X_I=0,  d(I) \ge N+1 \}.
\]
We can write a stratification
\[
\mathfrak{g}_N =\mathcal{V}_{1}\oplus...\oplus \mathcal{V}_{N},
\]
where
\[
\mathcal{V}_{k}=\mathbf{span} \{ U_I, d(I) =k\},\quad U_i=\Phi
(X_i).
\]
The canonical sublaplacian on $\mathbb{G}_N$ is
\[
\Delta_N =U_0+\frac{1}{2} \sum_{i=1}^d U_i^2.
\]
and, let us observe that the bracket generating condition is
satisfied at each point, so that due to H\"ormander's theorem,
$\Delta_N$ is hypoelliptic.

According to the Chow's theorem in Carnot groups (see for instance
Theorem 2.4 in \cite{Bau}), for $x_0 \in \mathbb{M}$, there exists
a unique smooth map
\[
\pi^N_{x_0}:\mathbb{G}_{N}  \rightarrow \mathbb{M}
\]
such that for any piecewise smooth path $x: [0,1] \rightarrow
\mathbb{R}^{d}$ and any smooth function $f: \mathbb{M} \rightarrow
\mathbb{R}$,
\[
\left[ \exp \left(  \sum_{I, d(I) \le N } \Lambda_I (x)_1 U_I
\right)( f \circ \pi^N_{x_0}) \right]( 1_{\mathbb{G}_N}) =\left[
\exp \left( \sum_{I, d(I) \le N } \Lambda_I (x)_1 V_I \right)f
\right] (x_0).
\]

Since
\[
Y_t=\exp \left(  \sum_{I, d(I) \le N } \Lambda_I (B)_t U_I \right)
(1_{\mathbb{G}_N})
\]
is seen to be the solution of the following stochastic
differential equation defined on $\mathbb{G}_{N}$
\[
Y_t =1_{\mathbb{G}_N}+\int_0^t Y_s \left( U_0 ds + \sum_{i=1}^d
U_i \circ dB^i_s \right).
\]
we get the following factorization of $\mathbf{Q}_t^N$ in
$\mathbb{G}_{N}$: For every $x_0 \in \mathbb{M}$ and every smooth
$f : \mathbb{M} \rightarrow \mathbb{R}$,
\[
(\mathbf{Q}^N_t f) (x_0)=e^{t \Delta_N} (f \circ \pi^N_{x_0})
(1_{\mathbb{G}_N}), \quad t \ge 0 .
\]
Since $\Delta_N$ is subelliptic, $e^{t \Delta_N}$ has a smooth
Schwartz kernel. But now, since $\mathcal{L}$ is elliptic at
$x_0$, the differential of $\pi^N_{x_0}$ has maximal rank at
$1_{\mathbb{G}_N}$. We get therefore the existence of
$\mathcal{O}_{x_0}$ and of a smooth
$q^N(x_0,\cdot):\mathbb{R}_{>0} \times
\mathcal{O}_{x_0}\rightarrow \mathbb{R}_{\ge 0}$ such that for
every every smooth $f :\mathbb{M} \rightarrow \mathbb{R} $ with
compact support included in $\mathcal{O}_{x_0}$,
\[
(\mathbf{Q}^N_t f)(x_0)=\int_\mathbb{M} q^N_t (x_0,y)f(y)dy.
\]

\

\textbf{Second step:}

\ For $t>0$, let us consider the operator $\Theta^N_t (x_0)$
defined on $\Gamma (\mathbb{M} , \mathcal{E})$ by the property
that for $\eta \in \Gamma (\mathbb{M} , \mathcal{E})$ and $ y \in
\mathcal{O}_{x_0}$,
\[
(\Theta^N_t (x_0) \eta)(y)=\mathbb{E} \left( \left[\exp \left(
\sum_{I, d(I) \le N } \Lambda_I (B)_t \nabla_{I} \right)\eta
\right] (x_0) \left| \exp \left( \sum_{I, d(I) \le N } \Lambda_I
(B)_t V_{I} \right)(x_0)=y\right) \right. .
\]
We claim that $\Theta^N_t (x_0)$ is actually a potential, that is
a smooth section of the bundle $\mathbf{End} (\mathcal{E})$. For
that, we have to show that for every smooth $f: \mathbb{M}
\rightarrow \mathbb{R}$ and every $\eta \in \Gamma (\mathbb{M} ,
\mathcal{E})$, $ y \in \mathcal{O}_{x_0}$,
\[
(\Theta_t (x_0) f\eta)(y)=f(y)(\Theta^N_t (x_0) \eta)(y).
\]
If $f$ is a smooth function on $\mathbb{M}$, we denote by
$\mathcal{M}_f$ the operator on $\Gamma (\mathbb{M} ,
\mathcal{E})$ that acts by multiplication by $f$. Due to the
Leibniz rule for connections, we have for any word $I$:
\[
[\nabla_I, \mathcal{M}_f] =\mathcal{M}_{V_I f}.
\]
Consequently,
\[
\left[ \sum_{I, d(I) \le N } \Lambda_I (B)_t \nabla_{I}
,\mathcal{M}_f \right]= \mathcal{M}_{\sum_{I, d(I) \le N }
\Lambda_I (B)_t V_{I}}.
\]
The above commutation property implies the following one:
\[
\exp \left( \sum_{I, d(I) \le N } \Lambda_I (B)_t \nabla_{I}
\right) \mathcal{M}_f=\mathcal{M}_{\exp \left( \sum_{I, d(I) \le N
} \Lambda_I (B)_t V_{I} \right) f}\exp \left( \sum_{I, d(I) \le N
} \Lambda_I (B)_t \nabla_{I} \right).
\]
Therefore,
\[
[\Theta^N_t(x_0), \mathcal{M}_f]=0,
\]
so that $\Theta^N_t(x_0) \in \Gamma(\mathbb{M},
\mathbf{End}(\mathcal{E}))$. We can now conclude with the
disintegration formula that for every $\eta \in \Gamma (
\mathbb{M}, \mathcal{E} ) $ with compact support included in
$\mathcal{O}_{x_0}$,
\[
(\mathbf{P}^N_t \eta)(x_0)=\int_\mathbb{M} p^N_t (x_0,y)\eta(y)dy,
\]
with
\[
p^N_t (x_0,\cdot)=q_t^N(x_0, \cdot) \Theta^N_t (x_0).
\]

\

\textbf{Final step:} \

Let us now turn to the proof of the pointwise estimate
\[
p_t (x_0,x_0)=p^N_t (x_0,x_0)  + O\left(
t^{\frac{N+1-d}{2}}\right), \quad t \rightarrow 0.
\]
Let $y \in \mathbb{M}$ sufficiently close to $x_0$. Since
$\mathcal{L}$ is elliptic at $x_0$, it is known that $p_t (x_0,y)$
admits a development
\begin{align}\label{devpt}
p_t (x_0,y)=\frac{e^{-\frac{d^2(x_0,y)}{2t}}}{(2\pi
t)^{d/2}}\left( \sum_{k=0}^N \Psi_k(x_0,y) t^k + t^{\frac{N+1}{2}}
\mathbf{R}_N (t,x_0,y) \right),
\end{align}
where the remainder term $\mathbf{R}_N (t,x_0,y)$ is bounded when
$t \rightarrow 0$, $\Psi_k (x_0, \cdot)$ is a section of
$\mathbf{End} (\mathcal{E})$ defined around $x_0$ and
$d(\cdot,\cdot)$ is the distance defined around $x_0$ by the
vector fields $V_1,...,V_d$. By using the fact that for every
smooth $f : \mathbb{M} \rightarrow \mathbb{R}$,
\[
(\mathbf{Q}^N_t f) (x_0)=e^{t \Delta_N} (f \circ \pi^N_{x_0})
(1_{\mathbb{G}_N}), \quad t \ge 0 ,
\]
and classical results for asymptotic development in small times of
subelliptic heat kernels (see for instance \cite{Ben}), we get for
$q^N_t (x_0,y)$ a development that is similar to (\ref{devpt}).
For $\Theta^N_t (x_0)$, the scaling property of Brownian motion
implies that we have a short-time asymptotics in powers
$t^{\frac{k}{2}}$, $k \in \mathbb{N}$. Since,
\[
p^N_t (x_0,\cdot)=q_t^N(x_0, \cdot) \Theta^N_t (x_0),
\]
we deduce that
\[
p^N_t (x_0,y)=\frac{e^{-\frac{d^2(x_0,y)}{2t}}}{(2\pi
t)^{d/2}}\left( \sum_{k=0}^N \tilde{\Psi}_k(x_0,y) t^k +
t^{\frac{N+1}{2}} \tilde{\mathbf{R}}_N (t,x_0,y) \right),
\]
where the remainder term $\tilde{\mathbf{R}}_N (t,x_0,y)$ is
bounded when $t \rightarrow 0$. With Theorem \ref{Parametrix}, we
obtain that $\Psi_k=\tilde{\Psi}_k$, $k=0,...,N$, and the required
estimate easily follows.
\end{proof}

If $I\in \{0,1,...,d\}^k$ is a word, we denote
\[
\mathcal{F}_I=\nabla_I-\nabla_{V_I} \in \Gamma
(\mathbb{M},\mathbf{End}(\mathcal{E})).
\]

\begin{corollary}
For $N \ge 1$, when $t \rightarrow 0$,
\[
p_t (x_0,x_0)=q^N_t (x_0)\mathbb{E} \left( \exp \left( \sum_{I,
d(I) \le N } \Lambda_I (B)_t \mathcal{F}_{I}(x_0) \right) \left|
\sum_{I, d(I) \le N } \Lambda_I (B)_t V_{I}(x_0)=0\right) \right.
+ O\left( t^{\frac{N+1-d}{2}}\right),
\]
where $q^N_t(x_0)$ is the density at $0$ of the random variable $
\sum_{I, d(I) \le N } \Lambda_I (B)_t V_{I}(x_0)$.
\end{corollary}

\section{The local index theorem for the Dirac operator on the
spin bundle}

The Atiyah-Singer index theorem for the Dirac operator on the spin
bundle as proved in \cite{At-Si}, is the following:

\begin{theorem}
Let $\mathbb{M}$ be a compact, $d$-dimensional spin manifold, with
$d$ even. Let $\mathbf{D}$ be the Dirac operator on the spin
bundle of $\mathbb{M}$. Then
\[
\mathbf{ind} (\mathbf{D} )=\left( \frac{1}{2i\pi}
\right)^{\frac{d}{2}} \int_\mathbb{M} [A(\mathbb{M})]_d,
\]
where $[A(\mathbb{M})]_d$ is the volume form on $\mathbb{M}$
obtained by taking the $d$-form piece of the $A$-genus
\[
A(\mathbb{M})=\det \left( \frac{ \Omega}{2 \sinh \frac{1}{2}
\Omega} \right)^{\frac{1}{2}},
\]
and $\Omega$ is the Riemannian curvature form defined in local
orthonormal frame $e_i$ with dual frame $e^*_i$ by
\[
\Omega=\frac{1}{2} \sum_{1 \le i,j \le d} R(e_i,e_j) e_i^* \wedge
e^*_j,
\]
with $R$, Riemannian curvature.
\end{theorem}

Before we turn to the proof. Let us first recall some linear
algebra constructions as can be found in Chapter 3 of
\cite{Ber-Ge-Ve}.

\

Let $V$ be an oriented $d$ dimensional Euclidean space. We assume
that the dimension $d$ is even. The Clifford algebra
$\mathbf{Cl}(V)$ over $V$ is the algebra
\[
\mathbf{T} (V)=\mathbb{R} \oplus V \oplus (V \otimes V) \oplus
\cdots
\]
quotiented by the relations
\begin{align}\label{relation-clifford}
u \otimes v +v \otimes u +2 \langle u,v \rangle 1 =0.
\end{align}
Let $e_1,...,e_d$ be an oriented basis of $V$. The family
\[
e_{i_1} ... e_{i_k}, \quad 0 \le k \le d, \quad 1 \le i_1 <...<i_k
\le d,
\]
forms a basis of $\mathbf{Cl}(V)$ that is therefore of dimension
$2^d$. In $\mathbf{T} (V)$ we can distinguish elements that are
even from elements that are odd. This leads to a decomposition:
\[
\mathbf{Cl}(V)=\mathbf{Cl}^-(V) \oplus \mathbf{Cl}^+(V),
\]
with $V \subset \mathbf{Cl}^-(V)$.

A Clifford module is a vector space $E$ over $\mathbb{R}$ (or
$\mathbb{C}$) that is also a $\mathbf{Cl}(V)$-module and that
admits a direct sum decomposition
\[
E=E^- \oplus E^+
\]
with
\[
\mathbf{Cl}^-(V) \cdot E^- \subset E^-, \quad \mathbf{Cl}^+(V)
\cdot E^+ \subset E^+.
\]
It can be shown that there is a unique Clifford module $S$, called
the spinor module over $V$ such that:
\[
\mathbf{End} (S) \simeq \mathbb{C} \otimes \mathbf{Cl}(V).
\]
In particular $\dim S=2^{\frac{d}{2}}$. There is therefore a
natural notion of supertrace on $\mathbf{Cl}(V)$ that is given by
\[
\mathbf{Str} \text{ } a =\mathbf{Tr}_{S^+} \text{
}a-\mathbf{Tr}_{S^-} \text{ }a,
\]
where $a \in \mathbf{Cl}(V)$ is seen as an element of
$\mathbf{End} (S)$. If
\[
a =\sum_{k=0}^d \sum_{1 \le i_1 <...<i_k \le d } a_{i_1 \cdots
i_k} e_{i_1} ... e_{i_k},
\]
then we have
\begin{align}\label{supertraceformula-clifford}
\mathbf{Str} \text{ } a=\left( \frac{2}{i} \right)^{\frac{d}{2}}
a_{1 \cdots d}.
\end{align}

If $\psi \in \mathfrak{so} (V)$, that is if $\psi : V \rightarrow
V$ is a skew-symmetric map, we define
\[
D\psi=\frac{1}{2}\sum_{1 \le i<j \le d} \langle \psi (e_i), e_j
\rangle e_i e_j \in \mathbf{Cl}(V),
\]
and observe that $D[\psi_1,\psi_2]=[D\psi_1,D\psi_2]$. The set
$\mathbf{Cl}^2(V)=D \mathfrak{so} (V)$ is therefore a Lie algebra.
The Lie group $\mathbf{Spin} (V)$ is the group obtained by
exponentiating $\mathbf{Cl}^2(V)$ inside the Clifford algebra
$\mathbf{Cl}(V)$; It is the two-fold universal covering of the
orthogonal group $\mathbf{SO}(V)$. It can also be described as the
set of $a \in \mathbf{Cl}(V)$ such that:
\[
a=v_1 ... v_{2k}, \quad 1 \le k \le \frac{d}{2},\quad v_i \in V,
\quad \parallel v_i \parallel=1.
\]

We now come back to differential geometry and carry the above
construction on the cotangent spaces of a spin manifold.

\

So, let $\mathbb{M}$ be a compact $d$-dimensional, Riemannian and
oriented manifold. We assume that $d$ is even. We furthermore
assume that $\mathbb{M}$ admits a spin structure: That is, there
exists a principal bundle on $\mathbb{M}$ with structure group
$\mathbf{Spin}(\mathbb{R}^d)$ such that the bundle charts are
compatible with the universal covering
$\mathbf{Spin}(\mathbb{R}^d) \rightarrow
\mathbf{SO}(\mathbb{R}^d)$. This bundle will be denoted
$\mathcal{SP} (\mathbb{M})$ and $\pi$ will denote the canonical
surjection. The spin bundle $\mathcal{S}$ over $\mathbb{M}$ is the
vector bundle such that for every $x \in \mathbb{M}$,
$\mathcal{S}_x$ is the spinor module over the cotangent space
$\mathbf{T}^*_x \mathbb{M}$. At each point $x$, there is therefore
a natural action of $\mathbf{Cl} (\mathbf{T}^*_x \mathbb{M})\simeq
\mathbf{End} (\mathcal{S}_x)$; this action will be denoted by
$\mathbf{c}$.

\

On $\mathcal{S}$, there is a canonical elliptic first-order
differential operator called the Dirac operator and denoted
$\mathbf{D}$. In a local orthonormal frame $e_i$, with dual frame
$e_i^*$, we have
\[
\mathbf{D}=\sum_{i} c(e^*_i) \nabla_{e_i},
\]
where $\nabla$ is the Levi-Civita connection. We have an analogue
of Weitzenb\"ock formula which is the celebrated Lichnerowicz
formula (see Theorem 3.52 in \cite{Ber-Ge-Ve}):
\[
\mathbf{D}^2=\Delta +\frac{s}{4},
\]
where $s$ is the scalar curvature of $\mathbb{M}$ and $\Delta$ is
given in a local orthonormal frame $e_i$ by
\[
\Delta=-\sum_{i=1}^d (\nabla_{e_i}\nabla_{e_i}- \nabla_{
\nabla_{e_i} e_i }).
\]

After these preliminaries, we can now turn to the proof of the
local index theorem for $\mathbf{D}$.

\

The first crucial step in all  heat equations approaches of index
theorems is the McKean-Singer type formula (see \cite{McK-Sin} and
Theorem 3.50 in \cite{Ber-Ge-Ve}):
\[
\mathbf{ind} (\mathbf{D})=\mathbf{Str}
(\mathbf{P}_t)=\int_{\mathbb{M}}\mathbf{Str}\text{ } p_t (x,x) dx
, \quad t >0,
\]
where $\mathbf{P}_t=e^{-\frac{1}{2} t \mathbf{D}^2}$, $p_t$ is the
corresponding Schwartz kernel, and $dx$ is the Riemannian volume
form.

By using the results of Section 3 we now show that we actually
have
\[
\lim_{t \rightarrow 0} \mathbf{Str}\text{ } p_t (x,x)dx=\left(
\frac{1}{2i\pi} \right)^{\frac{d}{2}}[A(\mathbb{M})]_d (x).
\]
This last statement is first due to Patodi \cite{Pat} and Gilkey
\cite{Gi} and implies the index theorem.

\

Let us  fix $x_0 \in \mathbb{M}$ once time for all in what
follows. Let $e_i$ be a synchronous local orthonormal frame
centered at $x_0$ with dual frame $e_i^*$. If needed, with a
cut-off function, we extend smoothly the vector field $e_i$ to be
zero outside a neigborhood of $x_0$. At the center $x_0$ of the
frame, we have:

\[
\Delta=-\sum_{i=1}^d \nabla_{e_i} \nabla_{e_i}
\]
and
\[
[\nabla_{e_i},\nabla_{e_j}]=R (e_i,e_j),
\]
where $R$ is the Riemannian curvature.

\ For $t >0$, let $\Theta_t (x_0) \in \mathbf{End}
(\mathcal{S}_{x_0}) \simeq \mathbf{Cl} (\mathbf{T}^*_{x_0}
\mathbb{M})$ be the Clifford element such that for every smooth
section $\eta$ of $\mathcal{S}$,
\[
\Theta_t (x_0) \eta(x_0)=\mathbb{E} \left( \left[\exp \left(
\sum_{I,d(I) \le d } \Lambda_I (B)_t \nabla_{I} \right)\eta
\right] (x_0) \left| \exp \left( \sum_{I, \mid I \mid \le d, 0
\notin I } \Lambda_{I} (B)_t e_{I} \right)(x_0)=x_0\right)\right.
,
\]
where in the above summation, we use the convention that
$\nabla_0$ is the multiplication operator by $-\frac{s}{8}$ and
$\nabla_i =\nabla_{e_i}, \text{ }1 \le i \le d$.
\begin{proposition}\label{spin-holonomy}
\[
\lim_{t \rightarrow 0} \frac{\mathbf{Str} \text{ } \Theta_t
(x_0)}{t^{d/2}}=\frac{1}{2^{d/2} (d/2)!}\mathbf{Str} \text{ }
\mathbb{E} \left(
 \left. \left(\sum_{1 \le i<j \le
d } D R (e_i,e_j) \int_0^1 B^i_s dB^j_s-B^j_s dB^i_s
\right)^{\frac{d}{2}} \right| B_1=0 \right)
\]
\end{proposition}

\begin{proof}

First, let us observe that due to the scaling property of Brownian
motion, for every smooth section $\eta$ of $\mathcal{S}$,
\[
\Theta_t (x_0) \eta(x_0)=\mathbb{E} \left( \left[\exp \left(
\sum_{I,d(I) \le d } t^{d(I)/2} \Lambda_I (B)_1 \nabla_{I}
\right)\eta \right] (x_0) \left| \exp \left( \sum_{I, \mid I \mid
\le d, 0 \notin I }t^{d(I)/2} \Lambda_{I} (B)_1 e_{I}
\right)(x_0)=x_0\right)\right.
\]

Let us now rewrite more explicitly $\Theta_t (x_0)$ as a Clifford
element.

If $1 \le i <j \le d$, we have at $x_0$
\[
[\nabla_{i},\nabla_{j}]=\mathbf{c} \left( D R (e_i,e_j) \right)
\]
with
\[
D R (e_i,e_j)=\frac{1}{2}\sum_{1 \le k<l \le d} \langle R(e_i,e_j)
e_k, e_l \rangle e^*_k e^*_l \in \mathbf{Cl}^2 (\mathbf{T}^*_{x_0}
\mathbb{M})
\]
Since the Levi-Civita connection is a Clifford connection, if $1
\le i <j <k \le d$, we have at $x_0$,
\begin{align*}
[\nabla_i,[\nabla_j,\nabla_k]]&=[\nabla_i,\nabla_{[e_j,e_k]} +
\mathbf{c} \left( D R
(e_j,e_k) \right)] \\
 & =\nabla_{[e_i,[e_j,e_k]]} + \mathbf{c} \left(DR (e_i,[e_j,e_k])+ \nabla_i D R
(e_j,e_k) \right).
\end{align*}
and we observe that $DR (e_i,[e_j,e_k])+ \nabla_i D R (e_j,e_k)
\in \mathbf{Cl}^2 (\mathbf{T}^*_{x_0} \mathbb{M})$.

 More generally, a recurrence procedure shows that if $1 \le
i_1 <...<i_k \le d$, then at $x_0$,
\[
\nabla_I-\nabla_{e_I}=c(\mathcal{F}_I),
\]
where $\mathcal{F}_I \in \mathbf{Cl}^2 (\mathbf{T}^*_{x_0}
\mathbb{M})$.

If $0 \in I$, then it is seen that, at $x_0$, $\nabla_I$ acts by
multiplication with a scalar. Therefore
\[
\Theta_t (x_0)=\mathbb{E} \left( \mathbf{c} \left(\exp \left(X_t+
\sum_{I,\mid I \mid \le d,0 \notin I } t^{\mid I \mid/2} \Lambda_I
(B)_1 \mathcal{F}_{I} \right) \right)  \left| \exp \left( \sum_{I,
\mid I \mid \le d, 0 \notin I }t^{\mid I \mid/2} \Lambda_{I} (B)_1
e_{I} \right)(x_0)=x_0\right)\right. ,
\]
where $X_t$ is scalar term such that $X_0=0$. We deduce
\[
\mathbf{Str} \text{ }\Theta_t (x_0)= \mathbb{E} \left(e^{X_t}
\mathbf{Str} \text{ }\exp \left(\sum_{I,\mid I \mid \le d,0 \notin
I } t^{\mid I \mid/2} \Lambda_I (B)_1 \mathcal{F}_{I} \right)
\left| \exp \left( \sum_{I, \mid I \mid \le d, 0 \notin I }t^{\mid
I \mid/2} \Lambda_{I} (B)_1 e_{I} \right)(x_0)=x_0\right)\right.
\]
and
\[
\mathbf{Str} \text{ }\Theta_t (x_0)\sim_{t \rightarrow 0}
\mathbb{E} \left( \mathbf{Str} \text{ } \left. \exp
\left(\sum_{I,\mid I \mid \le d,0 \notin I } t^{\mid I \mid/2}
\Lambda_I (B)_1 \mathcal{F}_{I} \right) \right| B_1=0\right)
\]
 But now, since $\mathcal{F}_I \in \mathbf{Cl}^2
(\mathbf{T}^*_{x_0} \mathbb{M})$, according to
(\ref{supertraceformula-clifford}), for any $k<\frac{d}{2}$,
\[
\mathbf{Str} \text{ } \left(\sum_{I,\mid I \mid \le d,0 \notin I }
t^{\mid I \mid/2} \Lambda_I (B)_1 \mathcal{F}_{I} \right)^k =0,
\]
and, when $t\rightarrow 0$,
\begin{align*}
 & \mathbb{E} \left( \mathbf{Str} \text{ } \left. \left(\sum_{I,\mid
I \mid \le d,0 \notin I } t^{\mid I \mid/2} \Lambda_I (B)_1
\mathcal{F}_{I} \right)^{\frac{d}{2}} \right| B_1=0 \right) \\
=& t^{d/2}\mathbb{E} \left( \mathbf{Str} \text{ } \left.
\left(\frac{1}{2}\sum_{1 \le i<j \le d } D R (e_i,e_j) \int_0^1
B^i_s dB^j_s-B^j_s dB^i_s  \right)^{\frac{d}{2}} \right| B_1=0
\right)+O(t^{\frac{d+1}{2}}).
\end{align*}

We conclude therefore

\[
 \lim_{t \rightarrow 0} \frac{\mathbf{Str} \text{ } \Theta_t
(x_0)}{t^{d/2}}=\mathbb{E} \left( \mathbf{Str} \text{ }
\frac{1}{(d/2)!} \left. \left(\frac{1}{2}\sum_{1 \le i<j \le d } D
R (e_i,e_j) \int_0^1 B^i_s dB^j_s-B^j_s dB^i_s
\right)^{\frac{d}{2}} \right| B_1=0 \right).
\]
\end{proof}

We can now obtain the required limit:

\begin{theorem}
\[
\lim_{t \rightarrow 0}\mathbf{Str} \text{ }p_t
(x_0,x_0)dx_0=\left( \frac{1}{2i\pi}
\right)^{\frac{d}{2}}[A(\mathbb{M})]_d (x_0).
\]
\end{theorem}

\begin{proof}
By Theorem \ref{main-estimation} and Proposition
\ref{spin-holonomy}, we get
\begin{align*}
& \lim_{t \rightarrow 0}\mathbf{Str} \text{ }p_t (x_0,x_0)dx_0 \\
=& \frac{1}{(4\pi)^{d/2}(d/2)!}\mathbf{Str} \text{ } \mathbb{E}
\left(
 \left. \left(\sum_{1 \le i<j \le
d } D R (e_i,e_j) \int_0^1 B^i_s dB^j_s-B^j_s dB^i_s
\right)^{\frac{d}{2}} \right| B_1=0 \right) e^*_1 \wedge ...
\wedge e^*_d .
\end{align*}
From (\ref{supertraceformula-clifford}) and from the expression of
$D R (e_i,e_j)$, the above expression is also equal to the
$d$-form piece of
\[
\frac{1}{(2i\pi)^{d/2}(d/2)!} \mathbb{E} \left(
 \left. \left(\sum_{1 \le i<j \le
d } \frac{1}{2} \Omega (e_i,e_j) \int_0^1 B^i_s dB^j_s-B^j_s
dB^i_s \right)^{\frac{d}{2}} \right| B_1=0 \right).
\]
This last expression is also the $d$-form piece of
\[
\frac{1}{(2i\pi)^{d/2}} \mathbb{E} \left(
 \left. \exp \left(\sum_{1 \le i<j \le
d } \frac{1}{2} \Omega (e_i,e_j) \int_0^1 B^i_s dB^j_s-B^j_s
dB^i_s \right) \right| B_1=0 \right),
\]
which turns out to be the $d$-form piece of the $A$-genus
\[
\frac{1}{(2i\pi)^{d/2}}A(\mathbb{M})=\frac{1}{(2i\pi)^{d/2}}\det
\left( \frac{ \Omega}{2 \sinh \frac{1}{2} \Omega}
\right)^{\frac{1}{2}},
\]
from L\'evy's area formula (see for instance Lemma 7.6.6 in
\cite{Hsu}).
\end{proof}

\end{document}